x^2-2x-15=0 by factoring, find the roots of the equation
Answer by checkley71(8403)

x^2-2x-15=0 Because the x^2 factor is =1 you need to find two numbers that satisfy the following criteria. n1*n2=-15 & n1+n2=-2.
First find all factors of -15 & they are:
-1&15
1&-15
-3&5
3&-5
now choose the set that when added =-2
-1+15=14 not an answer
1-15=-14 not an answer
-3+5=2 not an answer
3-5=-2 ANSWER! Now add thes two answers to the x term thus:
(x-5)(x+3)=0 Then set each term=0 (because for two numbers multiplied together &=0, one of them must be=0) and solve thus:
x-5=0
x=5 solution
x+3=0
x=-3 solution
---------------------------------------------
Hope this helps :)